# Bridged Posterior: Optimization, Profile Likelihood and a New Approach of Generalized Bayes

### v1

@@ Appendix B An Illustration of Least Favorable Submodel
and the $\lambda$-score function is $\dot{l}_{n}(\lambda,\zeta)=\sum_{i=1}^{n}\{x_{i}-x_{i}\exp(\lambda x_{i})Z(y_{i})\}/n$. Let $\widetilde{H}=L^{2}(\mathbb{R}_{+})$. Given a fixed function $\zeta_{0}$ and a bounded function $\delta\in\widetilde{H}$, we can define a path $\{\zeta_{\gamma}^{\delta}\in\mathcal{H}\}_{\gamma\in\mathbb{R}}$ by $\zeta_{\gamma}^{\delta}(y)=\{1+(\gamma-\lambda_{0})\delta(y)\}\zeta_{0}(y)$ for all $y\in\mathbb{R}_{+}$. It satisfies that $\zeta^{\delta}_{\gamma}\to\zeta_{0}$ in $\mathcal{H}$ when $\gamma\to\lambda_{0}$. Also, we define $Z_{\gamma}^{\delta}(y)=\int_{0}^{y}\{1+(\gamma-\lambda_{0})\delta(y)\}\zeta_{0}(y)\,\textup{d}y$ correspondingly. Now, plugging $\zeta^{\delta}_{\gamma}$ into $l_{n}(\lambda_{0},\zeta)$ as $\zeta$ and differentiating it at $\gamma=\lambda_{0}$, we get the $\zeta$-score function at $\zeta=\zeta_{0}$ in the direction of $\delta$ by
@@
Appendix C More Details of Proofs of Existing BvM Results
@@
In Bickel and Kleijn [2012], to get the LAN property for the marginal likelihood of $\lambda$, they first assume that a neighborhood of $z$ has enough prior mass, and the parameter space of $z$ has bounded Hellinger metric entropy (covering number). They assume that for $z$ outside a neighborhood of the fixed $z_{0}$, the Hellinger distances between the likelihood at $\lambda_{0}+h_{n}/\sqrt{n}$ and the one at $\lambda_{0}$ are uniformly (with respect to $z$) infinitesimal for every bounded $h_{n}$. Finally, assuming that the least favorable submodel exists, that is a submodel $l_{n}(\lambda,z_{\lambda})$ with parameters $(\lambda,z_{\lambda})$ satisfying $\mathscr{Q}\dot{l}_{n}(\lambda,z_{\lambda})=\dfrac{\partial l_{n}(\lambda,z_{\lambda})}{\partial\lambda}$ for all $\lambda$ in a neighborhood of $\lambda_{0}$, the conditional posterior distribution of $z$ can be shown to concentrate around the parameter of the least favorable submodel $z_{\lambda}$, that is the probability that the Hellinger distance between $z$ and $z_{\lambda}$ is greater than a positive number converges to zero. This leads to the marginal LAN property assuming that the full likelihood has LAN property in the direction of $\lambda$ when the $z$ is perturbed around the least favorable submodel, that is $z=z_{\lambda}+\zeta$ for all $\zeta$ in a neighborhood of 00.
@@
In Castillo and Rousseau [2015], let $h$ be the least favorable direction satisfying $\mathscr{P}\dot{l}_{n}(\lambda_{0},z_{0})=A_{\lambda_{0},z_{0}}h.$ Let $\lambda_{t}=\lambda-t\tilde{I}_{0}^{-1}/\sqrt{n},$ and $z_{t}=z+th\tilde{I}_{0}^{-1}/\sqrt{n}$. Assume over the direction of $h$, the log full likelihood has the LAN property with a remainder term, and the difference between the remainder terms at $(\lambda,z)$ and at $(\lambda_{t},z_{t})$ converges to zero. Further, suppose that the ratio between the integral of the likelihood under the prior at $(\lambda_{t},z_{t})$ and the one at $(\lambda,z)$ converge to $1$ for all $(\lambda,z)$. They proved the BvM theorem where the mean is $\lambda_{0}$ plus the first order term of the LAN expansion.
@@ Appendix D Data Augmentation for Latent Normal Model
For the latent normal model with binary observations, we follow Polson et al. [2013] and use the following data augmentation:
@@
Appendix E Simulation for Comparing the Posterior Variances from Latent Normal Model and Latent Quadratic Exponential Model

### v2

@@ Appendix B An Illustration of Least Favorable Submodel
and the $\lambda$-score function is $\dot{l}_{n}(\lambda,\zeta)=\sum_{i=1}^{n}\{x_{i}-x_{i}\exp(\lambda x_{i})Z(y_{i})\}/n$.
Let $\widetilde{H}=L^{2}(\mathbb{R}_{+})$. Given a fixed function $\zeta_{0}$ and a bounded function $\delta\in\widetilde{H}$, we can define a path $\{\zeta_{\gamma}^{\delta}\in\mathcal{H}\}_{\gamma\in\mathbb{R}}$ by $\zeta_{\gamma}^{\delta}(y)=\{1+(\gamma-\lambda_{0})\delta(y)\}\zeta_{0}(y)$ for all $y\in\mathbb{R}_{+}$. It satisfies that $\zeta^{\delta}_{\gamma}\to\zeta_{0}$ in $\mathcal{H}$ when $\gamma\to\lambda_{0}$. Also, we define $Z_{\gamma}^{\delta}(y)=\int_{0}^{y}\{1+(\gamma-\lambda_{0})\delta(y)\}\zeta_{0}(y)\,\textup{d}y$ correspondingly. Now, plugging $\zeta^{\delta}_{\gamma}$ into $l_{n}(\lambda_{0},\zeta)$ as $\zeta$ and differentiating it at $\gamma=\lambda_{0}$, we get the $\zeta$-score function at $\zeta=\zeta_{0}$ in the direction of $\delta$ by
@@
We conduct a numerical experiment for the Cox regression model. We generate the data with sample size $n=500$. We sample the covariate $(x_{1},\dots,x_{n})$ independently from a standard normal distribution. We create a baseline hazard function $\zeta_{0}$ with piecewise constant over 5 intervals, where hazard rates are assigned to each of these intervals. We set the ground-truth parameter of interest $\lambda_{0}$ as $0.8$. Then we sample the survival time $(y_{1},\dots,y_{n})$ from the hazard function $\exp(\lambda_{0}x_{i})\zeta_{0}(y_{i})$. For the bridged posterior based on profile likelihood, $g(\zeta,y;\lambda)=-L(y,\zeta;\lambda)$, we set N$(0,5^{2})$ prior on $\lambda$. For the canonical integrated posterior model, we set the same prior on $\lambda$ and Gamma(1,1)11(1,1)( 1 , 1 ) priors on hazard rates in each interval. We run the MCMC samplers for $10000$ iterations and discard the first $2000$ as burn-ins. Figure 9 shows that the posterior distributions for $\lambda$ under the two models are similar to each other.
(a) Posterior distribution of $\lambda$ from the full Bayesian model.
(b) Posterior distribution of $\lambda$ from the bridged posterior.
Figure 9: The posterior densities of the parameter $\lambda$ in Cox regression model under sample size 500.
Appendix C Comparison with Existing BvM Results on Semi-parametric Models
The second category of BvM results relate to canonical Bayesian methodology involving integrated posterior $\Pi(\lambda\mid y)=\int\Pi(\lambda,\textup{d}\zeta\mid y)$ over a non-deterministic $\zeta$. Bickel and Kleijn (2012) proved a BvM result for marginal posterior distribution of $\lambda$ using the LAN property for the marginal likelihood of $\lambda$, which has similar form with (11). On the other hand, they additionally assume that the marginal posterior probability of $\lambda$ inside the neighborhood $B_{M_{n}/\sqrt{n}}(\lambda_{0})$ converges to $1$ for every $M_{n}\to\infty$. This condition is similar to but arguably stronger than the last condition in Theorem 3. Castillo and Rousseau (2015) proved a BvM result on a functional of the parameters, under an essentially necessary no-bias condition which is related to both the likelihood and the prior specification of ($\lambda,\zeta$).
@@ Appendix C Comparison with Existing BvM Results on Semi-parametric Models
In Bickel and Kleijn (2012), to get the LAN property for the marginal likelihood of $\lambda$, they first assume that a neighborhood of $z$ has enough prior mass, and the parameter space of $z$ has bounded Hellinger metric entropy (covering number). They assume that for $z$ outside a neighborhood of the fixed $z_{0}$, the Hellinger distances between the likelihood at $\lambda_{0}+h_{n}/\sqrt{n}$ and the one at $\lambda_{0}$ are uniformly (with respect to $z$) infinitesimal for every bounded $h_{n}$. Finally, assuming that the least favorable submodel exists, that is a submodel $l_{n}(\lambda,z_{\lambda})$ with parameters $(\lambda,z_{\lambda})$ satisfying $\mathscr{Q}\dot{l}_{n}(\lambda,z_{\lambda})=\dfrac{\partial l_{n}(\lambda,z_{\lambda})}{\partial\lambda}$ for all $\lambda$ in a neighborhood of $\lambda_{0}$, the conditional posterior distribution of $z$ can be shown to concentrate around the parameter of the least favorable submodel $z_{\lambda}$, that is the probability that the Hellinger distance between $z$ and $z_{\lambda}$ is greater than a positive number converges to zero. This leads to the marginal LAN property assuming that the full likelihood has LAN property in the direction of $\lambda$ when the $z$ is perturbed around the least favorable submodel, that is $z=z_{\lambda}+\zeta$ for all $\zeta$ in a neighborhood of 00.
In Castillo and Rousseau (2015), let $h$ be the least favorable direction satisfying
@@
$\mathscr{P}\dot{l}_{n}(\lambda_{0},z_{0})=A_{\lambda_{0},z_{0}}h.$ Let $\lambda_{t}=\lambda-t\tilde{I}_{0}^{-1}/\sqrt{n},$ and $z_{t}=z+th\tilde{I}_{0}^{-1}/\sqrt{n}$. Assume over the direction of $h$, the log full likelihood has the LAN property with a remainder term, and the difference between the remainder terms at $(\lambda,z)$ and at $(\lambda_{t},z_{t})$ converges to zero. Further, suppose that the ratio between the integral of the likelihood under the prior at $(\lambda_{t},z_{t})$ and the one at $(\lambda,z)$ converge to $1$ for all $(\lambda,z)$. They proved the BvM theorem where the mean is $\lambda_{0}$ plus the first order term of the LAN expansion.
@@ Appendix D Data Augmentation for Latent Normal Model
For the latent normal model with binary observations, we follow Polson et al. (2013) and use the following data augmentation:
@@
Appendix E Additional Simulation Results
E.1 Results of Posterior Approximation for Latent Normal Model
We conduct simulated experiments to show results of posterior approximation for the latent normal model. We use two algorithms: the integrated nested Laplace approximations (INLA) and the variational inference with mean field approximation. For the INLA, we use inla function with default option from the R package INLA. For the variational inference, we use normal distribution with a diagonal covariance matrix as the variational distribution, and pyro package in python. Figure 10 shows that the INLA produces an distribution estimate similar to the exact posterior. On the other hand, the variational inference severely underestimates the value of $b$.
(a) Posterior distribution of $(b,\tau)$ from INLA.
(b) Posterior distribution of $(b,\tau)$ from variational inference.
Figure 10: The posterior distributions of the covariance kernel parameters from the latent normal model (with sample size 1000) using integrated nested Laplace approximations (Panel a) and variational inference (Panel b) .
E.2 Simulation for Comparing the Posterior Variances from Latent Normal Model and Latent Quadratic Exponential Model
@@ E.1 Results of Posterior Approximation for Latent Normal Model
E.3 Flow network problem with uncertainty on some capacity values
Consider a directed network $G=(V,E)$ with $V$ the set of nodes, and $E$ the set of uni-directed edges $(i\to j)$, each associated with a capacity value $c_{ij}>0$. We have observation of flow $y_{ij}\in[0,c_{ij}]$ on $(i\to j)\in E$, and $y_{ij}=0$ if $(i\to j)\not\in E$. Geography studies often impose a mechanistic model where the flow network is operating close to its maximum capacity when the network is congested, associated with the following linear program:
$g(\zeta;\lambda)=\sum_{j:\,(s\to j)\in E}\zeta_{sj}$
subject to $\sum_{j:\,(i\to j)\in E}\zeta_{ij}\,\,-\sum_{k:\,(k\to i)\in E}\zeta_{ji}=0,\quad\forall i\in V\setminus\{s,t\},$
$0\leq\zeta_{ij}\leq\lambda_{ij},\quad\forall(i\to j)\in E_{*},$
$0\leq\zeta_{ij}\leq c_{ij},\quad\forall(i\to j)\in E\setminus E_{*},$
where $s$ is the source node where there is a positive total flow entering the node, and $t$ is the sink node where there is a negative total flow corresponding to leaving the network. The equality constraints above correspond to flow conservation at each node that is neither source or sink. In practice, we often have a subset of edges $E_{*}$ where the capacity values $c_{i,j}$ are not constant but contain great uncertainty (such as roads that are prone to accidents), motivating for a statistical model where those $c_{ij}$ are replaced by parameters $\lambda_{ij}$. Taking consideration of the measurement error in $y_{ij}$, we have the following bridged posterior:
$L(y,z;\lambda)\pi_{0}(\lambda)\propto\Biggl{[}\prod_{k=1}^{n}(\sigma^{2})^{-|E|/2}\exp\biggl{\{}-\frac{\sum_{(i\to j)\in E}(y^{k}_{ij}-z_{ij})^{2}}{2\sigma^{2}}\biggr{\}}\Biggr{]}\exp(-\rho\sum_{(i\to j)\in E_{*}}\lambda_{ij})$
where $z={\arg\min}_{\zeta}g(\zeta;\lambda)$, and we assign an exponential prior for each $\lambda_{ij}$.
We use networkx to generate a directed network with $40$ nodes and $371$ edges, and make sure there is one source and one sink. We generate the capacities $c_{ij}$ independently from $\text{Uniform}(2,10)$. The ground-truth maximum flows $z^{0}_{ij}$ are computed using built-in function from networkx. We select $5$ edges to form the set $E_{*}$; the corresponding $\lambda_{ij}$’s are the low-dimensional parameters of interest. To simulate the observations, we generate $n=500$ samples $y_{ij}^{k}\sim\text{N}(z^{0}_{ij},1.0)$ for each edge $(i\to j)\in E$. For the prior, we set $\rho=0.2$, and we use an Inverse-Gamma$(2,5)$ prior on $\sigma_{2}^{2}$. We run the random walk Metropolis–Hastings algorithm for $10000$ iterations, discarding the first $2000$ as burn-in and thinning at $10$. Figure 12 show the good mixing performance of the posterior sampling algorithm, which takes $0.045$ seconds per iteration on a 4.0 GHz processor. Figure 11 shows the posterior densities of $\lambda_{ij}$’s for $(i\to j)\in E_{*}$.
Figure 11: Flow value posteriors of $\lambda_{ij}$ for each network edge in $E_{*}$ from the bridged posterior; vertical lines depict ground-truth values $c_{ij}$.
Figure 12: The autocorrelation functions for posterior Markov chains of $\lambda_{ij}$’s (by boxplots, left) and $\sigma^{2}$ (right).